OpenStudy (anonymous):

Find the indicated limits. lim x ->-2 x + 2/x^2 - 4

OpenStudy (anonymous):

@electrokid can u help?

OpenStudy (anonymous):

could you rewrite it using the equation editor

OpenStudy (anonymous):

yes hold on

OpenStudy (anonymous):

factor, and cancel then replace \(x\) by \(-2\)

OpenStudy (anonymous):

\[\lim_{x \rightarrow -2} x + 2/x^2 - 4 \]

OpenStudy (anonymous):

hmm you mean \[\lim_{x\to-2}\frac{x+2}{x^2-4}\] can we factor the denominator? \(a^2-b^2=(a+b)(a-b)\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\lim_{x\to-2}\frac{x+2}{(x+2)(x-2)}\]

OpenStudy (anonymous):

now, since \[x\to-2,\quad x+2\ne0\] and hence can be cancelled from numerator and denominator

OpenStudy (anonymous):

\[\lim_{x\to-2}\frac{\cancel{x+2}}{\cancel{(x+2)}(x-2)}=\lim_{x\to-2}\frac{1}{x-2}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now, you can substitute the value fr "x"

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

so its 1/-4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can only cancel the factors if they are not zero

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Hey can u help me with this one Find the indicated limit limx->infinity (e^2x - 1)/(x)

OpenStudy (anonymous):

are u there?

OpenStudy (anonymous):

are u there?